OpenStudy (anonymous):

8. How many units and in what direction will the function g(x) = (x + 5)2 - 9 translate if the parent function is f(x) = x2? A. 9 units to the left 5 units down B. 9 units to the left 5 units up C. 5 units to the left and 9 units down D. 5 units to the left and 9 units up

OpenStudy (mathstudent55):

We need to deal with a translation in x (horizontal translation) and a translation in y (vertical translation.) Let's start with the translation in x. When you replace x by x - k, the function moves k units along the x axis (right or left). If k is positive, it moves right. If k is negative, it moves left. For the translation in x (the horizontal translation), you are comparing \(x^2\) with \((x + 5)^2\). Write x + 5 as x - k. What is k?

OpenStudy (anonymous):

k isnegative @mathstudent55

OpenStudy (mathstudent55):

Correct. x + 5 = x - (-5) = x - k, so k = -5, so this means a translation of 5 units to the left.

OpenStudy (mathstudent55):

Now we need to deal with the vertical translation.

OpenStudy (mathstudent55):

Think of \(g(x) = (x + 5)^2 - 9\) as \(y = (x + 5)^2 - 9 \) Now add 9 to both sides: \(y + 9 = (x + 5)^2 \) A similar rule applies to y. If you replace y by y - h, the function has a translation of h units vertically. It translates up for positive h and down for negative h. Here y was replaced by y + 9. Now express y + 9 as y - h. What is h?

OpenStudy (anonymous):

nine is also negative @mathstudent55

OpenStudy (mathstudent55):

Correct. That means 9 units down.